# Overcoming the Generalization Limits of SLM Finetuning for Shape-Based Extraction of Datatype and Object Properties

### v1

@@ 1 Introduction
Relation Extraction (RE) is a core task in information extraction, which involves identifying entities in unstructured text and classifying semantic relationships between them. It plays a crucial role in populating and enriching knowledge graphs (KGs) by transforming textual data into structured relational triples. Recent advances in NLP, notably with language models (LMs), have significantly improved RE systems. These models now offer more context-aware and adaptable extraction capabilities with minimal task-specific supervision. To date, the primary focus of RE research has been on frequently occurring relations and well-defined entity types. Canonical datasets Zhang et al. (2017); Elsahar et al. (2018); Yao et al. (2019) were introduced to support this direction at scale, often employing distant supervision to align textual content with existing KGs. Such approaches have yielded proficient models in extracting various relational facts encompassing heterogeneous entity classes and relation types within a unified framework. Nonetheless, extracting less frequent and more complex relational patterns remains challenging for LMs.
Task definition: Following the formulation established in Ringwald et al. (2025), we propose learning RDF pattern extractors $\mathcal{M}_{Db}$ by finetuning Small LMs (SLMs). Training such models relies on a dual base $Db\subseteq$ $W\times{G}$, where $W$ is a set of texts and $G$ the set of associated KGs.
@@
In Ringwald et al. (2025) we focused on the extraction of datatype properties. In this paper, we extend the pattern-based RDF extraction to SHACL shapes that also contain object properties. Moreover, we highlight the critical challenge of ensuring balanced performance across all properties defined within a given SHACL shape, including under-represented properties, and we propose several solutions to address this issue. To summarize, this paper addresses the three following research questions:
@@
  • RQ3. How do such models perform over a long-tail distribution of properties, and how can this distribution be taken into account to better deal with under-represented properties?
@@ 2.1 Ontology-Driven Extraction
Several approaches proposed guiding RE with a predefined schema or ontology. They enable the extraction of relations not merely as independent triples, but within a contextual framework considering typed entities and the domain and range of the targeted relations. An example would be specifically targeting DBpedia relations involving entities belonging to the DBpedia class dbo:Company. An illustration of this paradigm is the SPIRES framework Caufield et al. (2024) which supports the construction of prompts based on schema-aware information retrieval. Among the various corpora developed for the RE task, Text2KGBench Mihindukulasooriya et al. (2023) stands out as one of the few datasets explicitly designed to support ontology-driven RE.
Positioning: Our previous work on RE from Wikipedia Ringwald et al. (2025) aligns with these works, using SHACL shapes to represent ontology constraints. This showed encouraging results when finetuning SLMs to create RDF pattern-based relation extractors. However, it was limited to extracting datatype properties. In this paper we extend our pattern-based RE approach to the extraction of both datatype and object properties.
@@ 2.2 Dealing with Under-represented Knowledge
Seq2Seq models finetuned on massive datasets, such as REBEL Huguet Cabot and Navigli (2021) and GenIE Josifoski et al. (2022), generally offer good performance but are not well-suited to deal with rare properties, as shown in He et al. (2025); Josifoski et al. (2023). The authors of SynthIE attempt to address this issue by generating synthetic data using a Large Language Model (LLM) as a data generator, which is subsequently used to train an SLM. Despite the good performance of these models on synthetic test data, their results do not reach the same level of performance on real-world, noisy data, such as those covered by REBEL. This limitation becomes obvious when evaluating SynthIE’s performance on the long-tail distribution of properties in the REBEL dataset.
@@
Text2KGBench Mihindukulasooriya et al. (2023) was initialized using two LLMs (Vicuna and Alapaca), and the results show the difficulty of extracting graphs described by an ontology. The main explanation is the models’ inability to handle numerous under-represented properties. In such cases, using the macro F1 metric is crucial to highlight the performance disparities across properties; unfortunately, this analysis is rarely conducted in RE. Similarly, Arzt et al. (2025) compares the performances of LMs on a relation classification task focusing on biographic data. It shows LM limitations on extracting under-represented relations, and highlights the importance of cross-dataset evaluation. Considering biographic data, Allen-Zhu and Li (2025) emphasises the importance of scaling laws in learning a specific piece of knowledge and shows that minimal exposure to a rare case is required for good performance, i.e., a minimum number of times a given piece of knowledge is represented in the training set. More precisely, the experiments show that 100 exposures during a training process are insufficient, and 1,000 exposures is proposed as a minimal condition to learn a fact. However, these conclusions were drawn in the context of pretraining models for memorization. In our context, we are finetuning models to retrieve facts from text. Among the numerous insights from this work, we also note the finding that data augmentation does not harm—and may even improve—model performance. This observation supports the potential of synthetic data augmentation strategies Ding et al. (2024), especially in low-resource scenarios.
@@
To address data imbalance, a standard solution in machine learning is using re-weighted loss functions to inversely balance each class by its importance in the training set Lin et al. (2017); Jiang et al. (2025); Xu et al. (2022). However, only little work investigates it in the context of LMs and RE.
@@
Positioning: Before we can generalize RDF pattern-based RE approaches to a broad set of SHACL shapes, we must first address the challenge of handling the long-tail distribution of properties with a single shape. To this end, we investigate stratification and re-weighted loss techniques, as well as several training data construction strategies—including scaling, augmentation, and ensuring sufficient exposure per property. The latter allows us to verify whether the “sufficient exposure” hypothesis proposed by Allen-Zhu and Li (2025) at the fact level also applies at the property level.
@@ 2.3 Contributions
(1) We extend the RDF pattern-based relation extraction in Ringwald et al. (2025) to both datatype and object properties.
@@ 3.1 Knowledge Distillation
In Ringwald et al. (2025), we considered the distillation of an initial dual base $\mathcal{K}$ consisting of the 2022.09 DBpedia datadump [^3] DBpedia data dump used which contains 6,109,994 Wikipedia abstracts and their DBpedia graphs:
@@
Here we extend the maximal shape defined in  Ringwald et al. (2025)444dbo:Person shape containing datatype and object prop. that considered only the following set of datatype properties:
@@
As done in Ringwald et al. (2025), we perform a data enrichment aiming to compensate for potentially missing datatype property values with the following set of rules [^5] For details Rules related to datatype prop.:
@@ 3.4 Training Details
All the trained models follow the configuration described in our previous work Ringwald et al. (2025). We used the BART-base (140M parameters) pre-trained model. We linearised the graphs following the “TurtleLight one line and factorised” Ringwald et al. (2025) syntax which demonstrated the best performance. The models were finetuned on a single Nvidia Tesla V100 GPU using the same configuration: an inverse square root scheduler with an initial learning rate of 0.00005, 1,000 steps of warmup, and configured with an early stop mode with patience of 5 steps. We used the same prompt to finetune the models: “\$entity_URI : \$Abstract” where \$Abstract is a Wikipedia abstract and \$entity_URI the URI of the corresponding entity in DBpedia. Each model was finetuned using a 10-fold cross-validation.
@@ 5.1 Data Augmentation Strategies
In the previous section, we showed that scaling and over-representation of rare properties can help improve model performance. However, the dbo:alias property remains particularly challenging. This property appears only 1,398 times in our distilled knowledge base (see Table 1) and has consistently been sampled below the sufficient exposure threshold of 1,000 examples, as emphasised by prior research Allen-Zhu and Li (2025). Such situations are common in low-resource scenarios. To address this, we propose to enrich our training data using three methods described in the next sections. Additional details of the data augmentation approaches in Sections 5.1.2 and 5.1.3 are given in our Github repository [^13] https://github.com/datalogism/Kastor/blob/main/doc/AugStrategies.md.
@@ 5.2 Results
. To evaluate the produced models, we reused the cross-evaluation set introduced in Section 4. Building on insights from Ringwald et al. (2025), we also extended our analysis with a corrected evaluation set: $d_{C}$. This dataset helps assess whether errors observed in $d_{R}$ stem from the discovery of new relevant facts (in the case of false positives), from omissions, or from noise in the KG (in the case of false negatives). We manually annotated 1,470 errors—comprising all false positives (FP) and false negatives (FN) produced during the evaluation of $d_{R}$ by the models $\mathcal{M}D1$, $\mathcal{M}D$\_, $\mathcal{M}D10$, but also $\mathcal{M}D10A_{all}$, $\mathcal{M}D10A_{KR0}$, $\mathcal{M}D10A_{KR1}$ and $\mathcal{M}D4SE_{all}$.

### v2

@@ 1 Introduction
Relation Extraction (RE) is a core task in information extraction, which involves identifying entities in unstructured text and classifying semantic relationships between them. It plays a crucial role in populating and enriching knowledge graphs (KGs) by transforming textual data into structured relational triples. Recent advances in NLP, notably with language models (LMs), have significantly improved RE systems. These models now offer more context-aware and adaptable extraction capabilities with minimal task-specific supervision. To date, the primary focus of RE research has been on frequently occurring relations and well-defined entity types. Canonical datasets zhang2017tacred; elsahar-etal-2018-rex; yao-etal-2019-docred were introduced to support this direction at scale, often employing distant supervision to align textual content with existing KGs. Such approaches have yielded proficient models in extracting various relational facts encompassing heterogeneous entity classes and relation types within a unified framework. Nonetheless, extracting less frequent and more complex relational patterns remains challenging for LMs.
Task definition: Following the formulation established in DBLP:conf/esws/RingwaldGFMA25, we propose learning RDF pattern extractors $\mathcal{M}_{Db}$ by finetuning Small LMs (SLMs). Training such models relies on a dual base $Db\subseteq$ $W\times{G}$, where $W$ is a set of texts and $G$ the set of associated KGs.
@@
In DBLP:conf/esws/RingwaldGFMA25 we focused on the extraction of datatype properties. In this paper, we extend the pattern-based RDF extraction to SHACL shapes that also contain object properties. Moreover, we highlight the critical challenge of ensuring balanced performance across all properties defined within a given SHACL shape, including under-represented properties, and we propose several solutions to address this issue. To summarize, this paper addresses the two following research questions:
@@ 2.1 Ontology-Driven Extraction
Several approaches proposed guiding RE with a predefined schema or ontology. They enable the extraction of relations not merely as independent triples, but within a contextual framework considering typed entities and the domain and range of the targeted relations. An example would be specifically targeting DBpedia relations involving entities belonging to the DBpedia class dbo:Company. An illustration of this paradigm is the SPIRES framework 10.1093/bioinformatics/btae104 which supports the construction of prompts based on schema-aware information retrieval. Among the various corpora developed for the RE task, Text2KGBench 10.1007/978-3-031-47243-5_14 stands out as one of the few datasets explicitly designed to support ontology-driven RE.
Positioning: Our previous work on RE from Wikipedia DBLP:conf/esws/RingwaldGFMA25 aligns with these works, using SHACL shapes to represent ontology constraints. This showed encouraging results when finetuning SLMs to create RDF pattern-based relation extractors. However, it was limited to extracting datatype properties. In this paper we extend our pattern-based RE approach to the extraction of both datatype and object properties.
@@ 2.2 Dealing with Under-represented Knowledge
Seq2Seq models finetuned on massive datasets, such as REBEL huguet-cabot-navigli-2021-rebel-relation and GenIE josifoski-etal-2022-genie, generally offer good performance but are not well-suited to deal with rare properties, as shown in he-etal-2025-language; josifoski-etal-2023-exploiting. The authors of SynthIE attempt to address this issue by generating synthetic data using a Large Language Model (LLM) as a data generator, which is subsequently used to train an SLM. Despite the good performance of these models on synthetic test data, their results do not reach the same level of performance on real-world, noisy data, such as those covered by REBEL. This limitation becomes obvious when evaluating SynthIE’s performance on the long-tail distribution of properties in the REBEL dataset.
@@
Text2KGBench 10.1007/978-3-031-47243-5_14 was initialized using two LLMs (Vicuna and Alapaca), and the results show the difficulty of extracting graphs described by an ontology. The main explanation is the models’ inability to handle numerous under-represented properties. In such cases, using the macro F1 metric is crucial to highlight the performance disparities across properties; unfortunately, this analysis is rarely conducted in RE. Similarly, arzt2025relationextractionpatternmatching compares the performances of LMs on a relation classification task focusing on biographic data. It shows LM limitations on extracting under-represented relations, and highlights the importance of cross-dataset evaluation. Considering biographic data, AL2024-knowledge3 emphasises the importance of scaling laws in learning a specific piece of knowledge and shows that minimal exposure to a rare case is required for good performance, i.e., a minimum number of times a given piece of knowledge is represented in the training set. More precisely, the experiments show that 100 exposures during a training process are insufficient, and 1,000 exposures is proposed as a minimal condition to learn a fact. However, these conclusions were drawn in the context of pretraining models for memorization. In our context, we are finetuning models to retrieve facts from text. Among the numerous insights from this work, we also note the finding that data augmentation does not harm—and may even improve—model performance. This observation supports the potential of synthetic data augmentation strategies ding2024dataaugmentationusinglarge, especially in low-resource scenarios.
@@
To address data imbalance, a standard solution in machine learning is using re-weighted loss functions to inversely balance each class by its importance in the training set 8237586; Jiang_Chan_Xue_Liu_Guo_2025; xu-etal-2022-towards-realistic. However, only little work investigates it in the context of LMs and RE.
@@
Positioning: Before we can generalize RDF pattern-based RE approaches to a broad set of SHACL shapes, we must first address the challenge of handling the long-tail distribution of properties with a single shape. To this end, we investigate stratification and re-weighted loss techniques, as well as several training data construction strategies—including scaling, augmentation, and ensuring sufficient exposure per property. The latter allows us to verify whether the “sufficient exposure” hypothesis proposed by AL2024-knowledge3 at the fact level also applies at the property level.
@@ 2.3 Contributions
(1) We extend the RDF pattern-based relation extraction in DBLP:conf/esws/RingwaldGFMA25 to both datatype and object properties.
@@ 3.1 Knowledge Distillation
In DBLP:conf/esws/RingwaldGFMA25, we considered the distillation of an initial dual base $\mathcal{K}$ consisting of the 2022.09 DBpedia datadump [^3] DBpedia data dump used which contains 6,109,994 Wikipedia abstracts and their DBpedia graphs:
@@
Here we extend the maximal shape defined in  DBLP:conf/esws/RingwaldGFMA25444dbo:Person shape containing datatype and object prop. that considered only the following set of datatype properties:
@@
As done in DBLP:conf/esws/RingwaldGFMA25, we perform a data enrichment aiming to compensate for potentially missing datatype property values with the following set of rules [^5] For details Rules related to datatype prop.:
@@ 3.4 Training Details
All the trained models follow the configuration described in our previous work DBLP:conf/esws/RingwaldGFMA25. We used the BART-base (140M parameters) pre-trained model. We linearised the graphs following the “TurtleLight one line and factorised” DBLP:conf/esws/RingwaldGFMA25 syntax which demonstrated the best performance. The models were finetuned on a single Nvidia Tesla V100 GPU using the same configuration: an inverse square root scheduler with an initial learning rate of 0.00005, 1,000 steps of warmup, and configured with an early stop mode with patience of 5 steps. We used the same prompt to finetune the models: “\$entity_URI : \$Abstract” where \$Abstract is a Wikipedia abstract and \$entity_URI the URI of the corresponding entity in DBpedia. Each model was finetuned using a 10-fold cross-validation.
@@ 5.1 Data Augmentation Strategies
In the previous section, we showed that scaling and over-representation of rare properties can help improve model performance. However, the dbo:alias property remains particularly challenging. This property appears only 1,398 times in our distilled knowledge base (see Table 1) and has consistently been sampled below the sufficient exposure threshold of 1,000 examples, as emphasised by prior research AL2024-knowledge3. Such situations are common in low-resource scenarios. To address this, we propose to enrich our training data using three methods described in the next sections. Additional details of the data augmentation approaches in Sections 5.1.2 and 5.1.3 are given in our Github repository [^13] https://github.com/datalogism/Kastor/blob/main/doc/AugStrategies.md.
@@ 5.2 Results
. To evaluate the produced models, we reused the cross-evaluation set introduced in Section 4. Building on insights from DBLP:conf/esws/RingwaldGFMA25, we also extended our analysis with a corrected evaluation set: $d_{C}$. This dataset helps assess whether errors observed in $d_{R}$ stem from the discovery of new relevant facts (in the case of false positives), from omissions, or from noise in the KG (in the case of false negatives). We manually annotated 1,470 errors—comprising all false positives (FP) and false negatives (FN) produced during the evaluation of $d_{R}$ by the models $\mathcal{M}D1$, $\mathcal{M}D$\_, $\mathcal{M}D10$, but also $\mathcal{M}D10A_{all}$, $\mathcal{M}D10A_{KR0}$, $\mathcal{M}D10A_{KR1}$ and $\mathcal{M}D4SE_{all}$.